# Swiss Army Knife: Synergizing BiAses in Knowledge from Vision Foundation Models for Multi-Task Learning

### v1

@@ Abstract
Vision Foundation Models (VFMs) have demonstrated outstanding performance on numerous downstream tasks. However, due to their inherent representation biases originating from different training paradigms, VFMs exhibit advantages and disadvantages across distinct vision tasks. Although amalgamating the strengths of multiple VFMs for downstream tasks is an intuitive strategy, effectively exploiting these biases remains a significant challenge. In this paper, we propose a novel and versatile “Swiss Army Knife” (SAK) solution, which adaptively distills knowledge from a committee of VFMs to enhance multi-task learning. Unlike existing methods that use a single backbone for knowledge transfer, our approach preserves the unique representation bias of each teacher by collaborating the lightweight Teacher-Specific Adapter Path modules with the Teacher-Agnostic Stem. Through dynamic selection and combination of representations with Mixture-of-Representations Routers, our SAK is capable of synergizing the complementary strengths of multiple VFMs. Extensive experiments show that our SAK remarkably outperforms prior state of the arts in multi-task learning by 10% on the NYUD-v2 benchmark, while also providing a flexible and robust framework that can readily accommodate more advanced model designs.
@@ 1 Introduction
Therefore, recent works (Ranzinger et al., 2024; Shang et al., 2024; Sariyildiz et al., 2024) propose more efficient frameworks by distilling multiple VFM teachers into a single student model, which can deliver competitive results on downstream benchmarks. Despite the progress, this many-to-one distillation risks eliminating the representation biases of the VFM teachers, potentially limiting the model’s ability to capitalize on their individual strengths for specific tasks. Zong et al. (2024) further point out that biased information from VFMs can lead to performance degradation when naively fused. Moreover, when matching one student to multiple teachers, reconciling diverse biases in shared parameters could induce optimization conflicts. Our pilot study in Table 2 shows that the student trained by many-to-one distillation does not consistently surpass the teachers in their respective proficient tasks.
@@ 2.1 Quantitative Analysis
To quantitatively analyze the representation biases, we evaluate the performance of the three VFMs directly transferred to each downstream task. We first freeze the models to generate image representations based on their pretrained knowledge, and then train a decoder head to produce final predictions for each task. DINOv2 and CLIP operate at a resolution of 512 on the downstream dataset, while SAM uses an input size of 1,024 as required by its pipeline. All feature maps are resized to $1/4$ of the input resolution before being passed to the head. To quantify the advantages of VFMs over the conventional ImageNet-pretrained ViT backbone, we calculate their relative improvement over ViT for each task.
@@ 2.2 Qualitative Analysis
On the other hand, SAM produces exceptional details in both tasks due to its high input resolution, as demonstrated in the red and cyan boxes. In the red box, a complex scene shows a foreground flower blending into the background, yet SAM accurately detects and labels the background in the segmentation mask. Notably, the background is not annotated in the ground truth, as such precise masking requires significant time and effort. We regard SAM’s high resolution as its representation bias, as it stems directly from the model’s training paradigm. However, SAM’s limitation lies in its semantic knowledge, particularly when integrating semantics from multiple objects. This makes it difficult to attain high-quality semantic segmentation results, even with highly precise masks, echoing the quantitative analysis.
@@ 2.3 Importance of Preserving Representation Biases
In summary, the inherent representation biases in VFMs result in their uneven performance across tasks, with no single model achieving the best results in all areas, as reflected in our quantitative analysis. This motivates the idea of combining multiple VFMs to achieve optimal performance in all tasks. Existing methods (Ranzinger et al., 2024; Shang et al., 2024; Sariyildiz et al., 2024) propose a solution by distilling multiple VFMs into a single student model. However, given that the student model is shared by several teachers, an important question naturally arises: Should we respect their individual representation biases when combining diverse VFMs?
To answer this question, we conduct a pilot study where a student model is distilled from three aforementioned VFMs, using only linear aligners to match the student’s features with those of the teachers, following the setup of the state-of-the-art method (Ranzinger et al., 2024). In this approach, the representation biases are not explicitly preserved during distillation, leading the student to learn a unified representation aimed at simultaneously matching all three teachers. We then evaluate its performance on downstream tasks with the same settings as in our quantitative analysis in Section 2.1. We compare the student without biases to an oracle derived from the teachers by selecting the best-performing teacher for each task, which represents the optimal performance of teachers.
@@ 3.2 Teacher-Agnostic Stem & Teacher-Specific Adapter Path
We utilize the standard adapter structure (Houlsby et al., 2019), which includes a down-projection layer ${\mathbf{W}}_{down}\in\mathbb{R}^{d\times r}$, a GELU non-linearity (Hendrycks & Gimpel, 2016), and an up-projection layer ${\mathbf{W}}_{up}\in\mathbb{R}^{r\times d}$, where $r\ll d$ is the reduced dimension. As in prior works (Chen et al., 2022; Mercea et al., 2024), we integrate a learnable scaling factor $\alpha$ and a residual connection from the input ${\bm{R}}_{in}\in\mathbb{R}^{n\times d}$ to the output ${\bm{R}}_{out}\in\mathbb{R}^{n\times d}$, which can be formulated as:
${\bm{R}}_{out}=\alpha{\mathbf{W}}_{up}\texttt{GELU}({\mathbf{W}}_{down}{\bm{R}}_{in})+{\bm{R}}_{in}.$ (3)
For a committee of $N$ VFM teachers, we assign a TSAP module with adapters $\{a_{l}^{i}\},l\in\{0,1,\dots,L\}$ to the $i$-th teacher. We then select four evenly distributed blocks from its outputs $\{{\bm{R}}_{l}^{i}\}$ to form multi-level representations $\{{\bm{R}}^{i}\}_{s}=\{{\bm{R}}^{i}_{l}\},l\in{\mathbb{L}}_{s}=\{L/4,L/2,3L/4,L\}$. Similarly, we have shared multi-level representations $\{{\bm{Z}}\}_{s}$ from TAS. Benefiting from the lightweight adapters, our distilled student model maintains original efficiency, as each TSAP module accounting for less than 5% of the TAS parameters. Consequently, our SAK framework is able to preserve the representation biases from the teachers without significant increases in computational cost, memory usage, or storage.
@@ 3.4 Training Paradigm
Our training paradigm contains two stages, with teacher parameters always frozen. In the first stage, we train the student model on the ImageNet-1k dataset (Deng et al., 2009; Russakovsky et al., 2015), focusing on aligning the outputs of the TSAP modules with their respective VFM teachers. ImageNet is chosen due to its diverse and extensive image samples, providing a strong basis for effective knowledge transfer. To maintain fairness—given that conventional ViT backbones are also pretrained on ImageNet—we opt not to use other larger datasets like those utilized in VFMs and RADIO (Ranzinger et al., 2024). Following previous findings (Ranzinger et al., 2024; Shang et al., 2024), we employ a combination of cosine distance and smooth-L1 losses for distillation. Let ${\bm{T}}_{l}^{i}$ be the $i$-th teacher’s representation at a selected level $l\in{\mathbb{L}}_{s}$, the overall distillation loss is defined as:
@@ 4.1 Experimental Setup
Datasets. We conduct experiments on two widely-used multi-task datasets: PASCAL-Context (Mottaghi et al., 2014) with five vision tasks and NYUD-v2 (Silberman et al., 2012) with four tasks. Details can be found in Appendix B.
@@
Baselines. To evaluate the effectiveness of our method, we consider three categories of baselines: (1) Single-task baseline, where individual models are trained for each task using the same ViT-initialized architecture, and multi-task baseline, where a shared encoder and task-specific heads are trained jointly. (2) Multi-teacher VFM distillation approaches, namely RADIO (Ranzinger et al., 2024) and Theia (Shang et al., 2024). We use their released models as encoder backbones, coupled with the same task heads as ours. (3) State-of-the-art MTL models, which involves complicated encoder or decoder designs. We assess the overall performance of each model with MTL Gain $\Delta_{m}$ by calculating the average relative difference across all tasks compared to the single-task baseline (Maninis et al., 2019).
@@ 4.2 Main Results
Figure 3 presents a comparison between our proposed SAK and representative baseline methods on both PASCAL-Context and NYUD-v2 datasets, with all methods using the ViT-B backbones. On PASCAL-Context, SAK greatly boosts the performance in Semseg and Parsing, achieving an overall improvement of 1.66% over the previous SOTA. On NYUD-v2, our method establishes a new milestone across all four tasks, increasing the MTL Gain metric from the previous best of 6.33% to 11.11%. We provide more comprehensive comparisons on both datasets using the ViT-L backbones in Tables 3 and 4. Our approach consistently outperforms previous methods, achieving the best results on 7 out of 9 tasks and both MTL Gain metrics. Notably, SAK significantly surpasses the SOTAs in MTL (BFCI (Zhang et al., 2023b), MLoRE (Yang et al., 2024d)) by nearly 10% on NYUD-v2, all while using fewer parameters.
@@ 4.3 In-depth Analysis
Ablation study. An ablation study is conducted to discern the individual contributions of the main components in SAK, namely TSAP and MoR Router, as outlined in Table 3. We consider two model variants: (1) a model without the TSAP and MoR Router modules (row 1), which corresponds to the student distilled naively regardless of representation biases, as studied in Table 2; (2) a model distilled with TSAP in the first stage but trained without MoR Routers in the second stage (row 2), where the biased representations from multiple VFMs are simply added together. Firstly, our results confirm that our proposed TSAP effectively preserves the representation biases from the teachers as indicated by a higher average similarity between the student and teachers. Additionally, we prove that a simple fusion of diverse biased knowledge does not lead to an overall improvement and may even fall behind compared to the student without biases. With the synergization of TSAP and MoR Router, our proposed SAK not only preserves and reproduces the representation biases after distillation but also optimally capitalizes on these biases to maximize multi-task performance.
@@
Impact of teacher selection. To investigate whether knowledge from all teachers can be effectively incorporated into the student model and how each teacher contributes to downstream tasks, we experiment on different combinations of VFM teachers in Table 4. When using a single teacher, SAK effectively learns the teacher’s representation bias, as the student distilled from DINOv2 or CLIP performs well in segmentation tasks, while SAM’s student is better in tasks requiring finer details. Combining DINOv2 and CLIP continues to improve segmentation tasks, potentially due to their complementary strengths in localized feature learning and semantic understanding. Including SAM further benefits all tasks, leading to the best overall results. We also visualize the gating weights learned by our proposed MoR Routers at the lowest and highest levels in Figure 4. At the lowest level, where VFM teachers share more general knowledge about the details, tasks tend to rely on the shared TAS and SAM’s bias. Conversely, the representation biases become more pronounced at higher levels; therefore, the teacher-specific representations are predominantly selected.
@@ 5 Related Work
Knowledge Distillation of Vision Foundation Models. As large-scale generalists, Vision Foundation Models (VFMs) show superior performance in various tasks with minimal tuning, such as CLIP (Radford et al., 2021) for vision-language tasks, DINOv2 (Oquab et al., 2024) for fine-grained recognition, and SAM (Kirillov et al., 2023) for promptable segmentation. To reduce their computational demands while preserving performance, knowledge distillation (Buciluǎ et al., 2006; Hinton et al., 2014) has been widely adopted in compressing VFMs (Vemulapalli et al., 2024; Sun et al., 2023; Yang et al., 2024a). More recently, multiple VFMs are distilled into a single student to combine their strengths: SAM-CLIP (Wang et al., 2024a) merges CLIP into SAM via continual learning and distillation. RADIO (Ranzinger et al., 2024) learns from CLIP, DINOv2, and SAM to enhance performance on downstream tasks. Theia (Shang et al., 2024) further incorporates Depth Anything (Yang et al., 2024b), showing advantages in robot learning. Different from the straightforward distillation in these methods, we adaptively transfer knowledge from multiple teachers while retaining the unique representation biases to maximize their strengths for multiple tasks.
@@ Appendix A Additional Analysis of Representation Biases
Based on our analysis in Section 2, we further delve into the representations generated by VFMs to understand how the representation biases lead to different outcomes across tasks. In Figure 6, we create a pair of manipulated images by selecting several cat images from the dataset and converting them into dogs while preserving the same shape and pose using ControlNet (Zhang et al., 2023c). We place various sizes of cats and dogs in a shared background image to compose complex and challenging scenes. For the segmentation task, this ensures that the predictions should differ only in labels, while the mask shapes remain the same. We visualize the representations by calculating the L2 norm of each patch token.
The red boxes highlight why CLIP struggles to detect small objects, such as the smallest cat and dog. In the representations, small objects exhibit a similar pattern to neighboring background tokens, indicating that CLIP assigns them the same level of importance as the background. Consequently, CLIP fails to segment these small objects and also struggles to generate precise masks for medium-sized objects, as shown in the cyan boxes. Conversely, DINOv2 performs better at detecting small objects by leveraging representations where these objects receive more attention than the background. However, its semantic understanding is limited in complex scenarios. For instance, DINOv2 confuses the heads of the cat and dog in the cyan boxes, even though this is the most distinguishable feature between the two animals. SAM displays different behaviors, with its representations concentrating more on the edges in the images instead of the entire objects, which explains its strength in capturing pixel-level details and weakness in high-level knowledge, as mentioned in our quantitative analysis. When examining the differences between the representations of the two images, CLIP appears more sensitive to changes in the main objects than DINOv2 and SAM, further underscoring its semantic focus. We also observe the phenomenon of register tokens (Darcet et al., 2024) in CLIP’s representation, where these artifacts correspond to areas the model deems low-informative background.
@@
CLIP is trained by aligning image and text embeddings of numerous image-text pairs through contrastive learning, which incorporates rich semantic knowledge from language domain that enhances image-level understanding. However, CLIP tends to prioritize prominent objects in an image while overlooking smaller ones, because (1) the alignment only considers the global class token (patch tokens are optimized implicitly), and (2) the image captions mainly focus on primary contents. This possibly accounts for its lower performance compared to DINOv2 in our quantitative experiments.
SAM’s representation bias can be explained by two reasons. First, SAM is supervised by the promptable segmentation task, considering solely the object masks and ignoring their semantic labels. Second, SAM operates with an input resolution of 1,024 in both training and inference, producing nearly four times more patches than other models. Consequently, SAM is adept at capturing pixel-level details and detecting object edges but poor at perceiving the semantics of multiple objects.

### v2

@@ Abstract
Vision Foundation Models (VFMs) have demonstrated outstanding performance on numerous downstream tasks. However, due to their inherent representation biases originating from different training paradigms, VFMs exhibit advantages and disadvantages across distinct vision tasks. Although amalgamating the strengths of multiple VFMs for downstream tasks is an intuitive strategy, effectively exploiting these biases remains a significant challenge. In this paper, we propose a novel and versatile “Swiss Army Knife” (SAK) solution, which adaptively distills knowledge from a committee of VFMs to enhance multi-task learning. Unlike existing methods that use a single backbone for knowledge transfer, our approach preserves the unique representation bias of each teacher by collaborating the lightweight Teacher-Specific Adapter Path modules with the Teacher-Agnostic Stem. Through dynamic selection and combination of representations with Mixture-of-Representations Routers, our SAK is capable of synergizing the complementary strengths of multiple VFMs. Extensive experiments show that our SAK remarkably outperforms prior state of the arts in multi-task learning by 10% on the NYUD-v2 benchmark, while also providing a flexible and robust framework that can readily accommodate more advanced model designs. Project page: https://innovator-zero.github.io/SAK/.
@@ 1 Introduction
Therefore, recent works (Ranzinger et al., 2024b; Shang et al., 2024; Sariyildiz et al., 2024) propose more efficient frameworks by distilling multiple VFM teachers into a single student model, which can deliver competitive results on downstream benchmarks. Despite the progress, this many-to-one distillation risks eliminating the representation biases of the VFM teachers, potentially limiting the model’s ability to capitalize on their individual strengths for specific tasks. Zong et al. (2024) further point out that biased information from VFMs can lead to performance degradation when naively fused. Moreover, when matching one student to multiple teachers, reconciling diverse biases in shared parameters could induce optimization conflicts. Our pilot study in Table 2 shows that the student trained by many-to-one distillation does not consistently surpass the teachers in their respective proficient tasks.
@@ 2.1 Quantitative Analysis
To quantitatively analyze the representation biases, we evaluate the performance of the three VFMs directly transferred to each downstream task. We first freeze the models to generate image representations based on their pretrained knowledge, and then train a decoder head to produce final predictions for each task. DINOv2 and CLIP operate at a resolution of 512 on the downstream dataset, while SAM uses an input size of 1,024 as required by its pipeline. All feature maps are resized to $1/4$ of the output resolution before being passed to the head. To quantify the advantages of VFMs over the conventional ImageNet-pretrained ViT backbone, we calculate their relative improvement over ViT for each task.
@@ 2.2 Qualitative Analysis
On the other hand, SAM produces exceptional details in both tasks due to its high input resolution, as demonstrated in the red and cyan boxes. In the red box, a complex scene shows a foreground flower blending into the background, yet SAM accurately detects and labels the background in the segmentation mask. Notably, the background is not annotated in the ground truth, as such precise masking requires significant time and effort. We regard SAM’s high resolution as its representation bias, as it stems directly from the model’s training paradigm. However, SAM’s limitation lies in its semantic knowledge, particularly when integrating semantics from multiple objects. This makes it difficult to attain high-quality semantic segmentation results, even with highly precise masks, echoing the quantitative analysis. We provide additional analysis and discussions in Appendix A.
@@ 2.3 Importance of Preserving Representation Biases
In summary, the inherent representation biases in VFMs result in their uneven performance across tasks, with no single model achieving the best results in all areas, as reflected in our quantitative analysis. This motivates the idea of combining multiple VFMs to achieve optimal performance in all tasks. Existing methods (Ranzinger et al., 2024b; Shang et al., 2024; Sariyildiz et al., 2024) propose a solution by distilling multiple VFMs into a single student model. However, given that the student model is shared by several teachers, an important question naturally arises: Should we respect their individual representation biases when combining diverse VFMs?
To answer this question, we conduct a pilot study where a student model is distilled from three aforementioned VFMs, using only linear aligners to match the student’s features with those of the teachers, following the setup of the state-of-the-art method (Ranzinger et al., 2024b). In this approach, the representation biases are not explicitly preserved during distillation, leading the student to learn a unified representation aimed at simultaneously matching all three teachers. We then evaluate its performance on downstream tasks with the same settings as in quantitative analysis in Section 2.1. We compare the student without biases to an oracle derived from the teachers by selecting the best-performing teacher for each task, which represents the optimal performance of teachers.
@@ 3.2 Teacher-Agnostic Stem & Teacher-Specific Adapter Path
We utilize the standard adapter structure (Houlsby et al., 2019), which includes a down-projection layer ${\mathbf{W}}_{\text{down}}\in\mathbb{R}^{d\times r}$, a GELU non-linearity (Hendrycks & Gimpel, 2016), and an up-projection layer ${\mathbf{W}}_{\text{up}}\in\mathbb{R}^{r\times d}$, where $r\ll d$ is the reduced dimension. As in prior works (Chen et al., 2022; Mercea et al., 2024), we integrate a learnable scaling factor $\alpha$ and a residual connection from the input ${\bm{R}}_{\text{in}}\in\mathbb{R}^{n\times d}$ to the output ${\bm{R}}_{\text{out}}\in\mathbb{R}^{n\times d}$, which can be formulated as:
${\bm{R}}_{\text{out}}=\alpha\texttt{GELU}({\bm{R}}_{\text{in}}{\mathbf{W}}_{\text{down}}){\mathbf{W}}_{\text{up}}+{\bm{R}}_{\text{in}}.$ (3)
For a committee of $N$ VFM teachers, we assign a TSAP module with adapters $\{a_{l}^{i}\},l\in\{0,1,\dots,L\}$ to the $i$-th teacher. We then select four evenly distributed blocks from its outputs $\{{\bm{R}}_{l}^{i}\}$ to form multi-level representations $\{{\bm{R}}^{i}\}_{s}=\{{\bm{R}}^{i}_{l}\},l\in{\mathbb{L}}_{s}=\{L/4,L/2,3L/4,L\}$. Similarly, we have shared multi-level representations $\{{\bm{Z}}\}_{s}$ from TAS. Benefiting from the lightweight adapters, our distilled student model maintains original efficiency, as each TSAP module accounting for less than 5% of the TAS parameters. Consequently, our SAK framework is able to preserve the representation biases from the teachers without significant increases in computational cost, memory usage, or storage, as shown in detailed discussions in Appendix D.1.
@@ 3.4 Training Paradigm
Our training paradigm contains two stages, with teacher parameters always frozen. In the first stage, we train the student model on the ImageNet-1k dataset (Deng et al., 2009; Russakovsky et al., 2015), focusing on aligning the outputs of the TSAP modules with their respective VFM teachers. ImageNet is chosen due to its diverse and extensive image samples, providing a strong basis for effective knowledge transfer. To maintain fairness—given that conventional ViT backbones are also pretrained on ImageNet—we opt not to use other larger datasets like those utilized in VFMs and RADIO (Ranzinger et al., 2024b). Following previous findings (Ranzinger et al., 2024b; Shang et al., 2024), we employ a combination of cosine distance and smooth-L1 losses for distillation. Let ${\bm{T}}_{l}^{i}$ be the $i$-th teacher’s representation at a selected level $l\in{\mathbb{L}}_{s}$, the overall distillation loss is:
@@ 4.1 Experimental Setup
Datasets. We conduct experiments on two widely-used multi-task datasets: PASCAL-Context (Mottaghi et al., 2014) with five vision tasks and NYUD-v2 (Silberman et al., 2012) with four tasks. Details can be found in Appendix B.2.
@@
Baselines. To evaluate the effectiveness of our method, we consider three categories of baselines: (1) Single-task baseline, where individual models are trained for each task using the same ViT-initialized architecture, and multi-task baseline, where a shared encoder and task-specific heads are trained jointly. (2) Multi-teacher VFM distillation approaches, namely RADIO (Ranzinger et al., 2024b) and Theia (Shang et al., 2024). We use their released models as encoder backbones, coupled with the same task heads as ours. (3) State-of-the-art MTL models, which involves complicated encoder or decoder designs. We assess the overall performance of each model with MTL Gain $\Delta_{m}$ by calculating the average relative difference across all tasks compared to the single-task baseline (Maninis et al., 2019).
@@ 4.2 Main Results
Figure 3 presents a comparison between our proposed SAK and representative baseline methods on both PASCAL-Context and NYUD-v2 datasets, with all methods using the ViT-B backbones. On PASCAL-Context, SAK greatly boosts the performance in Semseg and Parsing, achieving an overall improvement of 1.66% over the previous SOTA. On NYUD-v2, our method establishes a new milestone across all four tasks, increasing the MTL Gain metric from the previous best of 6.33% to 11.11%. We provide more comprehensive comparisons on both datasets using the ViT-L backbones in Tables 3 and 4, and ViT-S/Swin-S backbones in Appendix C. Our approach consistently outperforms previous methods, achieving the best results on 7 out of 9 tasks and both MTL Gain metrics. Notably, SAK significantly surpasses the SOTAs in MTL (BFCI (Zhang et al., 2023b), MLoRE (Yang et al., 2024d)) by nearly 10% on NYUD-v2, all while using fewer parameters.
@@ 4.3 In-depth Analysis
Ablation study. An ablation study is conducted to discern the individual contributions of the main components in SAK, namely TSAP and MoR Router, as outlined in Table 3. We consider two model variants: (1) a model without the TSAP and MoR Router modules (row 1), which corresponds to the student distilled naively regardless of representation biases, as studied in Table 2; (2) a model distilled with TSAP in the first stage but trained without MoR Routers in the second stage (row 2), where the biased representations from multiple VFMs are simply added together. Firstly, our results confirm that our proposed TSAP effectively preserves the representation biases from the teachers as indicated by a higher average similarity between the student and teachers. Additionally, we prove that a simple fusion of diverse biased knowledge does not lead to an overall improvement and may even fall behind compared to the student without biases. With the synergization of TSAP and MoR Router, our proposed SAK not only preserves and reproduces the representation biases after distillation but also optimally capitalizes on these biases to maximize multi-task performance. The upper-bound results of teacher amalgamation are presented in Appendix C.
@@
Impact of teacher selection. To investigate whether knowledge from all teachers can be effectively incorporated into the student model and how each teacher contributes to downstream tasks, we experiment on different combinations of VFM teachers in Table 4. When using a single teacher, SAK effectively learns the teacher’s representation bias, as the student distilled from DINOv2 or CLIP performs well in segmentation tasks, while SAM’s student is better in tasks requiring finer details. Combining DINOv2 and CLIP continues to improve segmentation tasks, potentially due to their complementary strengths in localized feature learning and semantic understanding. Including SAM further benefits all tasks, leading to the best overall results. We also visualize the gating weights learned by our proposed MoR Routers at the lowest and highest levels in Figure 4. At the lowest level, where VFM teachers share more general knowledge about the details, tasks tend to rely on the shared TAS and SAM’s bias. Conversely, the representation biases become more pronounced at higher levels; therefore, the teacher-specific representations are predominantly selected. Further analysis is provided in Appendix D.3 and D.4.
@@ 5 Related Work
Knowledge Distillation of Vision Foundation Models. As large-scale generalists, Vision Foundation Models (VFMs) show superior performance in various tasks with minimal tuning, such as CLIP (Radford et al., 2021) for vision-language tasks, DINOv2 (Oquab et al., 2024) for fine-grained recognition, and SAM (Kirillov et al., 2023) for promptable segmentation. To reduce their computational demands while preserving performance, knowledge distillation (Buciluǎ et al., 2006; Hinton et al., 2014) has been widely adopted in compressing VFMs (Vemulapalli et al., 2024; Sun et al., 2023; Yang et al., 2024a). More recently, multiple VFMs are distilled into a single student to combine their strengths: SAM-CLIP (Wang et al., 2024a) merges CLIP into SAM via continual learning and distillation. RADIO (Ranzinger et al., 2024b) learns from CLIP, DINOv2, and SAM to enhance performance on downstream tasks. Theia (Shang et al., 2024) further incorporates Depth Anything (Yang et al., 2024b), showing advantages in robot learning. Different from the straightforward distillation in these methods, we adaptively transfer knowledge from multiple teachers while retaining the unique representation biases to maximize their strengths for multiple tasks.
@@ 6 Conclusion
Acknowledgments
This work was supported in part by NSF Grant 2106825, NIFA Award 2020-67021-32799, the Toyota Research Institute, the IBM-Illinois Discovery Accelerator Institute, the Amazon-Illinois Center on AI for Interactive Conversational Experiences, Snap Inc., and the Jump ARCHES endowment through the Health Care Engineering Systems Center at Illinois and the OSF Foundation. This work used computational resources, including the NCSA Delta and DeltaAI supercomputers through allocations CIS230012 and CIS240428 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, as well as the TACC Frontera supercomputer and Amazon Web Services (AWS) through the National Artificial Intelligence Research Resource (NAIRR) Pilot.
Ethics Statement
Our research adheres to high ethical standards in machine learning and computer vision, ensuring transparency, reproducibility, and fairness in all experiments. While our approach shows promising results, it shares common limitations with other foundation models, particularly regarding data usage. We utilize publicly available datasets in compliance with relevant legal and ethical guidelines, emphasizing proper data handling during selection and pre-processing.
Reproducibility Statement
We ensure reproducibility by elaborating implementation details, which includes details on models, datasets, and training setups in Appendix B. Our code and models are publicly available at https://github.com/innovator-zero/SAK.
@@ Appendix A Additional Analysis of Representation Biases
Based on our analysis in Section 2, we further delve into the representations generated by VFMs to understand how the representation biases lead to different outcomes across tasks. In Figure 6, we create a pair of manipulated images by selecting several cat images from the dataset and converting them into dogs while preserving the same shape and pose using ControlNet (Zhang et al., 2023c). We place various sizes of cats and dogs in a shared background image to compose complex and challenging scenes. For the segmentation task, this ensures that the predictions should differ only in labels, while the mask shapes remain the same. We visualize the representations by calculating the L2 norm of each patch token and their representation differences.
The red boxes highlight why CLIP struggles to detect small objects, such as the smallest cat and dog. In the representations, small objects exhibit a pattern similar to neighboring background tokens, indicating that CLIP assigns them the same level of importance as the background. Consequently, CLIP fails to segment these small objects and also struggles to generate precise masks for medium-sized objects, as shown in the cyan boxes. Conversely, DINOv2 performs better at detecting small objects by leveraging representations where these objects receive more attention than the background. However, its semantic understanding is limited in complex scenarios. For instance, DINOv2 confuses the heads of the cat and dog in the cyan boxes, even though this is the most distinguishable feature between the two animals. SAM displays different behaviors, with its representations concentrating more on the edges in the images instead of the entire objects, which explains its strength in capturing pixel-level details and weakness in high-level knowledge, as mentioned in our quantitative analysis. When examining the differences between the representations of the two images, CLIP appears more sensitive to changes in the main objects than DINOv2 and SAM, further underscoring its semantic focus. We also observe the phenomenon of register tokens (Darcet et al., 2024) in CLIP’s representation, where these artifacts correspond to areas the model deems low-informative background.
@@
  • CLIP is trained by aligning the image and text embeddings of numerous image-text pairs through contrastive learning, which incorporates rich semantic knowledge from language domain that enhances image-level understanding. However, CLIP tends to prioritize prominent objects in an image while overlooking smaller ones, because (1) the alignment only considers the global class token (patch tokens are optimized implicitly), and (2) the image captions mainly focus on primary contents. This possibly accounts for its lower performance compared to DINOv2 in our quantitative experiments.
  • SAM’s representation bias can be explained by two reasons. First, SAM is supervised by the promptable segmentation task, considering solely the object masks and ignoring their semantic labels. Second, SAM operates with an input resolution of 1,024 in both training and inference, producing nearly four times more patches than other models. Consequently, SAM is adept at capturing pixel-level details and detecting object edges, but is poor at perceiving the semantics of multiple objects.